# The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxies

### v1

@@ 2.2.2 Dwarf galaxy simulations
The low-metallicity ($Z=0.01$ $\mathrm{Z}_{\odot}$) dwarf galaxy initial conditions are adopted from Lahén et al. (2023), based on the dwarf galaxy models of Hu et al. (2016). We use the compact initial condition with virial mass of $4\times 10^{10}$ $\mathrm{M}_{\odot}$, including a $4\times 10^{7}$ $\mathrm{M}_{\odot}$ gas disk and a $2\times 10^{7}$ $\mathrm{M}_{\odot}$ old stellar disk with disk scale lengths of 0.73 kpc. The initial mass-resolution is 4 $\mathrm{M}_{\odot}$ for gas and disk stars and $\sim 6800$ $\mathrm{M}_{\odot}$ for dark matter. The gravitational force-softening length is set to 62 pc for dark matter, 0.1 pc for gas and 0.1 pc for pre-existing disk star particles. For newly formed stars we use $\epsilon_{*}$ of 0.1 pc or 0.01 pc (see Section 2.3 for details) except when for stars that are within ketju-regions in the ketju-simulations. The hydrodynamical simulations are named as hyd_KETJU when ketju is enabled and as hyd_SPHGAL when ketju is disabled.
@@ 2.4 Identification of star clusters
In the hydrodynamical dwarf galaxy simulations, we identify bound star clusters using friend-of-friends and subfind (Springel et al., 2001; Dolag et al., 2009), which are structure finding algorithms included in gadget-3. In order to follow the long-term evolution of clusters especially in the ketju-runs, we use a linking length of 1 pc to recover the extended distribution of stars. The subfind-algorithm has been modified to account for the non-softened gravitational potential within the ketju-regions in the runs where ketju is enabled. We require at least 50 bound stars to be counted as a cluster, and exclude clusters with a mean stellar age younger than 1 Myr when making direct comparisons to observations where deeply embedded, extremely young clusters are often excluded.
@@ 3.1 Lagrangian radii
The clusters in the simulations with regularisation start compact and their centres contract even further during the first few Myr while mass-segregation proceeds. In response, the outer regions expand, up to a factor of 10 and 100 for the 50% and 90% radii, respectively. The inner regions of individual nb_10k_BIFROST-runs and nb_10k_KETJU-runs go through cycles of gradual contraction (see Fig. 12 in Appendix The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxies) that is halted by binary formation and rapid expansion of $r_{5\%}$. The expansion is then finished through the ejection of massive star(s) from the centre. However, this is not visible after the first core collapse cycle (around $\sim 1$ Myr in the bifrost run) when the cluster size is averaged over 10 runs. This is because each random simulation realisation has a different time-scale for the collapse-cycle depending on the masses of the most massive stars still present in the cluster. Overall, the mean 5% Lagrangian radius increases after the first few Myr.
The simulations that either integrate all close interactions with ketju (nb_1k_KETJU_m0_e001) or include a larger fraction of the cluster stars in the ketju integration (nb_1k_KETJU_m3_e01) evolve almost identically to the direct N-body runs. The results of the other simulations that use ketju but only include stars above a mass threshold as the ketju region centres (KETJU_m3, KETJU_m8) are inbetween the direct N-body results and the fully softened runs. This signifies that the dynamical processes important for the internal evolution of the clusters, such as mass-segregation and core-collapse, are captured significantly better in all runs with ketju compared to fully softened runs. Comparing nb_10k_KETJU_m3_e001 and nb_10k_KETJU_m8_e001, we see that the dynamical evolution of the clusters is not very sensitive to the exact number of ketju-regions. In other words, even only accounting for the collisional evolution of a few per cent of the stars in the cluster results in significantly more realistic cluster evolution.
@@ 4.3 SN clustering
In Fig. 6 we show the ISM densities around young stars at formation, around photoionising young stars ($<1$ Myr) and around stars that have exploded as SN. To inspect how the accurate stellar dynamics impacts clustering of stellar feedback, we also show separately feedback events that occur in bound star clusters (bottom panel) recovered in both simulations (hyd_KETJU_m3_e01 and hyd_SPHGAL_m3_e01). While photoionisation and star formation occur at slightly lower densities in the ketju-run, the SN environmental densities are quite similar in both runs. This manifestly shows the impact of effective early stellar feedback in regulating the gaseous densities in star-forming regions. The only significant difference in the two runs is the number of SNe in clusters, which is consistent with Fig. 5 where we saw a lower recycling efficiency of wind and SN-material. The fraction of SNe in clusters drops from $68\%$ in the sphgal-simulation to $26\%$ in the ketju-run. The effect of lowered SN-clustering on the galaxy outflows was however very minor as shown in Fig. 5, with a difference of only 10% in the total gas mass that escapes the galactic disk between the two simulations.
@@ 4.4 Evolution of the star cluster population
The CMF in the ketju-runs differ from the fully softened simulations both at young and at old cluster ages. Firstly, the ketju-run has a shallower power-law shape at high cluster masses already at young cluster age. Observationally, the power-law slope of the CMF is often seen to be close to $-2$, thus the ketju simulations seem to be in better agreement with the observations. However, the observed cluster samples are often more massive ($>10^{3-4}$ $\mathrm{M}_{\odot}$), have higher metallicity or are older, and are rarely complete at such low masses as considered here. We do not therefore attempt a direct comparison here.
@@ 4.4.1 Expansion of young compact clusters
In Fig. 8 we show the effective radii reff (2D in x–y plane) of clusters in three age bins of 1–5 Myr, 5–10 Myr and 10–100 Myr. The age ranges have been selected to enable direct comparison with observed star clusters in dwarf galaxies. We compare our cluster sizes to cluster data from the LEGUS survey (Brown & Gnedin, 2021), here limited to low-mass galaxies with $M_{*}<10^{9}$ $\mathrm{M}_{\odot}$, and clusters observed in the LMC and SMC from Hunter et al. (2003) and Gatto et al. (2021). These three samples were selected as they cover the full age and mass range considered here and provide an estimate of the half-mass or half-light radius, instead of e.g. a core-radius or a 90% light radius as is often the case in Magellanic cloud cluster data. Fig. 8 shows also the best-fit relation from Brown & Gnedin (2021) and this relation scaled down by a factor of 8. The latter illustrates the initial cluster sizes expected from N-body simulations that start with compact configurations and end on the observed size-mass relation (Banerjee & Kroupa, 2017; Arca Sedda et al., 2024).
@@ 4.5.3 Gas expulsion
We further inspect the intersection of embedded and exposed clusters in the hyd_KETJU_m3_e01 ketju-run by showing the time of the birth of the first photoionising star and the first SN event within 20 pc of the cluster centre of mass superimposed on the mass-size evolution sequence from Fig. 10. The embedded phase comes to an end when the average stellar age is a couple Myr [^8] As opposed to the total stellar age spread or the total lifetime of the star cluster, which can be larger. and often coincident with the ignition of the first photoionizing star. The first SNe occur only Myrs later, and do not contribute to the gas removal of the host cluster. We note that in the simulations of the present study, the shortest lifetimes of the SN-progenitor stars with masses up to 40 $\mathrm{M}_{\odot}$ are $\sim 5$ Myr in the absence of stars in the pair-instability mass range. In a more intensely star-forming galaxy (e.g. Lahén et al. 2024a), the onset of SNe could be hastened through the formation of more massive star clusters that are able to host more massive stars with lifetimes as short as a couple Myrs.
The small sizes and the gas expulsion induced by photoionisation in our simulated young clusters therefore seem to be in agreement with observed star clusters, and a picture emerges wherein the clusters evolve rapidly toward lower central densities during and after gas removal. The driver of gas removal is photoionising radiation, well before SNe in the vicinity of the clusters come into play. This result is in agreement with previous simulations that have shown that a sufficiently luminous source can significantly disrupt the gas cloud before the SNe occur (Dale et al., 2013; Geen et al., 2016; Kim et al., 2018; Guszejnov et al., 2022), and our results extend these previous works with a model that includes individually resolved radiation sources in a full galactic context. The present results are also in qualitative agreement with previous analysis of the simulations in the griffin-project on starburst systems (Fotopoulou et al. 2024) where we have shown that even the most massive cold star-forming ISM-clouds ($>10^{5}$ $\mathrm{M}_{\odot}$) can be destroyed already after the first SN. The pre-processing by early stellar feedback is thus a crucial driver of evolution of clustered star-forming regions across environments. The drastic difference in the size-evolution of the collisional and softened clusters in Fig. 10 however highlights the fact that the full impact of gas removal on the evolution of star clusters can only be addressed in hydrodynamical simulations that account for their collisional nature.
@@ 5 Summary and Conclusions
In this study we discuss clusters less massive than 1000 $\mathrm{M}_{\odot}$, whereas gas retention might be more efficient in more massive clusters as indicated in our previous work on star-by-star starburst environments in Lahén et al. (2024b). Investigation of more massive clusters will be left for a future study, where we will apply the new hydrodynamical+N-body method in a low-metallicity starburst. An implementation for stellar interactions and mergers would enable us to study the collisional growth of very massive stars in dense star clusters (Portegies Zwart et al., 2004). Such dense cluster-forming regions may also have given rise to globular clusters in the early Universe (Adamo et al., 2024). The rapid (Myr-scale) collisional growth of massive stars in these dense clusters may provide a channel for enhanced chemical enrichment (Gieles et al., 2018) and black hole formation in the mass range of $\sim 10^{3}$ $\mathrm{M}_{\odot}$$10^{4}$ $\mathrm{M}_{\odot}$ (Fujii et al., 2024; Rantala et al., 2024). With this simulation methodology we may begin to address the origin of the chemical enhancements (Bunker et al., 2023) and the seeds of the supermassive black holes (Bogdán et al., 2024) recently uncovered with JWST at unexpectedly high redshifts.
@@ Acknowledgements
N.L. and T.N. gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SUPERMUC-NG at Leibniz Supercomputing Centre (www.lrz.de) under project numbers pn49qi and pn72bu. TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS". P.H.J. acknowledges the support by the European Research Council via ERC Consolidator Grant KETJU (no. 818930) and the support of the Academy of Finland grant 339127. The computations were carried out at SuperMUC-NG hosted by the LRZ and the COBRA and FREYA clusters hosted by The Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.
This research made use of python packages scipy (Virtanen et al., 2020), numpy (Harris et al., 2020), matplotlib (Hunter, 2007), pygad (Röttgers et al., 2020) and h5py (Collette, 2013).
@@
Appendix A Central 5% Lagrangian radii of individual clusters

### v2

@@ 2.2.2 Dwarf galaxy simulations
The low-metallicity ($Z=0.01$ $\mathrm{Z}_{\odot}$) dwarf galaxy initial conditions are adopted from Lahén et al. (2023), based on the dwarf galaxy models of Hu et al. (2016). We use the compact initial condition with virial mass of $4\times 10^{10}$ $\mathrm{M}_{\odot}$, including a $4\times 10^{7}$ $\mathrm{M}_{\odot}$ gas disk and a $2\times 10^{7}$ $\mathrm{M}_{\odot}$ old stellar disk with disk scale lengths of 0.73 kpc. The initial mass-resolution is 4 $\mathrm{M}_{\odot}$ for gas and disk stars and $\sim 6800$ $\mathrm{M}_{\odot}$ for dark matter. The gravitational force-softening length is set to 62 pc for dark matter, 0.1 pc for gas and 0.1 pc for pre-existing disk star particles. The smallest SPH smoothing lengths containing the 100 nearest gas particles are $\sim 0.3$ pc, corresponding to densities of $10^{5}$ cm-3 at the typical star formation threshold. For newly formed stars we use $\epsilon_{*}$ of 0.1 pc or 0.01 pc (see Section 2.3 for details) except when for stars that are within ketju-regions in the ketju-simulations. The hydrodynamical simulations are named as hyd_KETJU when ketju is enabled and as hyd_SPHGAL when ketju is disabled.
@@ 2.4 Identification of star clusters
In the hydrodynamical dwarf galaxy simulations, we identify bound star clusters using friends-of-friends and subfind (Springel et al., 2001; Dolag et al., 2009), which are structure finding algorithms included in gadget-3. In order to follow the long-term evolution of clusters especially in the ketju-runs, we use a linking length of 1 pc to recover the extended distribution of stars. The subfind-algorithm has been modified to account for the non-softened gravitational potential within the ketju-regions in the runs where ketju is enabled. We require at least 50 bound stars to be counted as a cluster, and exclude clusters with a mean stellar age younger than 1 Myr when making direct comparisons to observations where deeply embedded, extremely young clusters are often excluded. The cluster finding is performed for each snapshot over a timespan of 500 Myr in 1 Myr intervals.
@@ 3.1 Lagrangian radii
The clusters in the simulations with regularisation start compact and their centres contract even further during the first few Myr while mass-segregation proceeds. In response, the outer regions expand, up to a factor of 10 and 100 for the 50% and 90% radii, respectively. The inner regions of individual nb_10k_BIFROST-runs and nb_10k_KETJU-runs go through cycles of gradual contraction (see Fig. 12 in Appendix The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxies) that is halted by binary formation and rapid expansion of $r_{5\%}$. The expansion is then finished through the ejection of massive star(s) from the centre. However, this is not visible after the first core collapse cycle (around $\sim 1$ Myr in the bifrost run) when the cluster size is averaged over 10 runs. This is because each random simulation realisation has a different time-scale for the collapse-cycle depending on the masses of the most massive stars still present in the cluster. Overall, the mean 5% Lagrangian radius increases after the first few Myr, as the clusters have too low masses to undergo further core collapse (see Fig. 12 and e.g. Heggie & Ramamani 1989).
The simulations that either integrate all close interactions with ketju (nb_1k_KETJU_m0_e001) or include a larger fraction of the cluster stars in the ketju integration (nb_1k_KETJU_m3_e01) evolve almost identically to the direct N-body runs. The results of the other simulations that use ketju but only include stars above a mass threshold as the ketju region centres (KETJU_m3, KETJU_m8) are inbetween the direct N-body results and the fully softened runs. This signifies that the dynamical processes important for the internal evolution of the clusters, such as mass-segregation and core collapse, are captured significantly better in all runs with ketju compared to fully softened runs. Comparing nb_10k_KETJU_m3_e001 and nb_10k_KETJU_m8_e001, we see that the dynamical evolution of the clusters is not very sensitive to the exact number of ketju-regions. In other words, even only accounting for the collisional evolution of a few per cent of the stars in the cluster results in significantly more realistic cluster evolution.
@@ 4.3 SN clustering
In Fig. 6 we show the ISM densities around young stars at formation, around photoionising young stars ($<1$ Myr) and around stars that have exploded as SN. To inspect how the accurate stellar dynamics impacts clustering of stellar feedback, we also show separately feedback events that occur in bound star clusters (bottom panel) recovered in both simulations (hyd_KETJU_m3_e01 and hyd_SPHGAL_m3_e01). While photoionisation and star formation occur at slightly lower densities in the ketju-run, the SN environmental densities are quite similar in both runs. This manifestly shows the impact of effective early stellar feedback in regulating the gaseous densities in star-forming regions. The only significant difference in the two runs is the number of SNe in clusters, which is consistent with Fig. 5 where we saw a lower recycling efficiency of wind and SN-material. The number of SNe in bound clusters reduces from 993 to 509, out of 1428 and 1273 SNe in total, respectively. The fraction of SNe in clusters therefore drops from $70\%$ in the sphgal-simulation to $40\%$ in the ketju-run. The effect of lowered SN-clustering on the galaxy outflows was however very minor as shown in Fig. 5, with a difference of only 10% in the total gas mass that escapes the galactic disk between the two simulations.
@@ 4.4 Evolution of the star cluster population
The CMF in the ketju-runs differ from the fully softened simulations both at young and at old cluster ages. Firstly, the ketju-run has a shallower power-law shape at high cluster masses already at young cluster age. Observationally, the power-law slope of the CMF is often seen to be close to $-2$, thus the ketju simulations seem to be in better agreement with the observations. However, the observed cluster samples are often more massive ($>10^{3-4}$ $\mathrm{M}_{\odot}$), have higher metallicity or are older, and are rarely complete at such low masses as considered here. We do not therefore attempt a direct comparison here. Other simulations of star cluster formation (e.g. Hislop et al. 2022; Garcia et al. 2023; Andersson et al. 2024) have previously found that the shape of the CMF and mass of the the most massive cluster can also depend strongly on the stellar feedback and star formation prescriptions.
@@ 4.4.1 Expansion of young compact clusters
In Fig. 8 we show the effective radii reff (2D in x–y plane) of clusters in three age bins of 1–5 Myr, 5–10 Myr and 10–100 Myr. The age ranges have been selected to enable direct comparison with observed star clusters in dwarf galaxies. We compare our cluster sizes to cluster data from the LEGUS survey (Brown & Gnedin, 2021), here limited to low-mass galaxies with $M_{*}<10^{9}$ $\mathrm{M}_{\odot}$, and clusters observed in the LMC and SMC from Hunter et al. (2003) and Gatto et al. (2021). The radii used in Hunter et al. (2003) are from Bica et al. (1999). These three samples were selected as they cover the full age and mass range considered here and provide an estimate of the half-mass or half-light radius, instead of e.g. a core-radius or a 90% light radius as is often the case in Magellanic cloud cluster data. Fig. 8 shows also the best-fit relation from Brown & Gnedin (2021) and this relation scaled down by a factor of 8. The latter illustrates the initial cluster sizes expected from N-body simulations that start with compact configurations and end on the observed size-mass relation (Banerjee & Kroupa, 2017; Arca Sedda et al., 2024).
@@ 4.5.3 Gas expulsion
We further inspect the intersection of embedded and exposed clusters in the hyd_KETJU_m3_e01 ketju-run in Fig. 11 by showing the time of the birth of the first photoionising star and the first SN event within 20 pc of the cluster centre of mass superimposed on the mass-size evolution sequence from Fig. 10. The embedded phase comes to an end when the average stellar age is a couple Myr [^8] As opposed to the total stellar age spread or the total lifetime of the star cluster, which can be larger. and often coincident with the ignition of the first photoionizing star. The first SNe occur only Myrs later, and do not contribute to the gas removal of the host cluster. We note that in the simulations of the present study, the shortest lifetimes of the SN-progenitor stars with masses up to 40 $\mathrm{M}_{\odot}$ are $\sim 5$ Myr in the absence of stars in the pair-instability mass range. In a more intensely star-forming galaxy (e.g. Lahén et al. 2024), the onset of SNe could be hastened through the formation of more massive star clusters that are able to host more massive stars with lifetimes as short as a couple Myrs.
The small sizes and the gas expulsion induced by photoionisation in our simulated young clusters therefore seem to be in agreement with observed star clusters, and a picture emerges wherein the clusters evolve rapidly toward lower central densities during and after gas removal. The driver of gas removal is photoionising radiation, well before SNe in the vicinity of the clusters come into play. This result is in agreement with previous simulations that have shown that a sufficiently luminous source can significantly disrupt the gas cloud before the SNe occur (Dale et al., 2013; Geen et al., 2016; Kim et al., 2018; He et al., 2019, 2020; Fukushima et al., 2020; Guszejnov et al., 2022), and our results extend these previous works with a model that includes individually resolved radiation sources in a full galactic context. The present results are also in qualitative agreement with previous analysis of the simulations in the griffin-project on starburst systems (Fotopoulou et al. 2024) where we have shown that even the most massive cold star-forming ISM-clouds ($>10^{5}$ $\mathrm{M}_{\odot}$) can be destroyed already after the first SN. The pre-processing by early stellar feedback is thus a crucial driver of evolution of clustered star-forming regions across environments, though full gas removal by photoionisation might be only possible in relatively low-mass clusters. The drastic difference in the size-evolution of the collisional and softened clusters in Fig. 10 however highlights the fact that the full impact of gas dispersal and removal on the evolution of star clusters can only be addressed in hydrodynamical simulations that account for their collisional nature. The internal evolution of more massive clusters with a prolonged embedded phase will be addressed in a future study.
@@ 5 Summary and Conclusions
In this study we discuss clusters less massive than 1000 $\mathrm{M}_{\odot}$. The star formation time scale is longer (He et al., 2019; Lahén et al., 2020) and the gas retention may be more efficient in more massive clusters as indicated in our previous work on star-by-star starburst environments in Lahén et al. (2024). In more massive clusters, where the gas can be retained in the cluster centre for a longer period of time, stellar winds and SN feedback may become more important for chemical enrichment and gas removal. Investigation of more massive clusters will be left for a future study, where we will apply the new hydrodynamical+N-body method in a low-metallicity starburst. An implementation for stellar interactions and mergers would enable us to study the collisional growth of very massive stars in dense star clusters (Portegies Zwart et al., 2004). Such dense cluster-forming regions may also have given rise to globular clusters in the early Universe (Adamo et al., 2024), which may have been important sources of ionising radiation (He et al., 2020). The rapid (Myr-scale) collisional growth of massive stars in these dense clusters may provide a channel for enhanced chemical enrichment (Gieles et al., 2018) and black hole formation in the mass range of $\sim 10^{3}$ $\mathrm{M}_{\odot}$$10^{4}$ $\mathrm{M}_{\odot}$ (Fujii et al., 2024; Rantala et al., 2024). With this simulation methodology we may begin to address the origin of the chemical enhancements (Bunker et al., 2023) and the seeds of the supermassive black holes (Bogdán et al., 2024) recently uncovered with JWST at unexpectedly high redshifts.
@@ Acknowledgements
NL thanks Deidre Hunter and Bruce Elmegreen for providing the data reported in Hunter et al. (2003). NL and TN gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SUPERMUC-NG at Leibniz Supercomputing Centre (LRZ, www.lrz.de) under project numbers pn49qi and pn72bu. TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS". PHJ acknowledges the support by the European Research Council via ERC Consolidator Grant KETJU (no. 818930) and the support of the Academy of Finland grant 339127. The computations were carried out at SuperMUC-NG hosted by the LRZ and the COBRA and FREYA clusters hosted by The Max Planck Computing and Data Facility (MPCDF) in Garching, Germany.
This research made use of python packages scipy (Virtanen et al., 2020), numpy (Harris et al., 2020), matplotlib (Hunter, 2007), pygad (Röttgers et al., 2020), h5py (Collette, 2013) and astropy [^9] http://www.astropy.org (Astropy Collaboration et al., 2022).
@@
Appendix A Central 5 per cent Lagrangian radii of individual clusters